OpenStudy (anonymous):

factoring trinomials in type ax^2+bx+c. factor completely. if it is a prime state it. 6x^3-4x^2-10x

OpenStudy (anonymous):

what are the steps in factoring it?

OpenStudy (anonymous):

6x^3-4x^2-10x = 2x(3x^2 - 2x - 5) = 2x[3x^2 - 5x + 3x -5] = 2x[x(3x - 5) +1(3x - 5)] = 2x[(x + 1)(3x - 5)] = 2x(x + 1)(3x - 5)

OpenStudy (anonymous):

First we take out the common factor 2x For the polynomial in the bracket a=3, b= -2 and c= -5 Then we calculate a*c = 3*(-5) = -15 Next we find two factors of -15, say X and Y, such that X*Y = a*c i.e. -15 and X+Y = b = -2 With a little work we find the combination of -5 and 3 rest is self-explanatory....... Hope that explains it for you................☺

OpenStudy (anonymous):

yes it does i understand it pretty clearly now. I just did not understand the factoring part itself. Thank you very much for helping.